x + 4y = -10
2x – 3y = 13

Answer by mananth(16949)

1 x + 4 y = -10 .............1
2 x -3 y = 13 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 4
3 x 12 y = -30
8 x -12 y = 52
Add the two equations
11 x = 22
/ 11
x = 2
plug value of x in (1)
1 x + 4 y = -10
2 + 4 y = -10
4 y = -10 -2
4 y = -12
y = -3
(2,3) is the solution